OpenStudy (mendicant_bias):

I found a particularly ugly indefinite integral that I can't seem to figure out how to evaluate (technique used is supposed to be u-substitution). Anybody care to take a gander?

OpenStudy (mendicant_bias):

Ugh, I'm not able to pull up the equation editor.

OpenStudy (mendicant_bias):

The integrand is: sqrt((x^3 + 3)/x^11). Supposed to figure out how to deal with that via u-sub. @hartnn , any clue?

OpenStudy (mendicant_bias):

Correction, -3, but that shouldn't really matter in this case.

OpenStudy (mendicant_bias):

I think I just maybe, heh...lol, keep going, but I think I'm figuring something out

OpenStudy (mendicant_bias):

Nope, nevermind, don't got it.

ganeshie8 (ganeshie8):

pull out x^4 from the denominator

ganeshie8 (ganeshie8):

then u may see wat to sub

OpenStudy (mendicant_bias):

One sec, trying to figure out if I can get the eqn editor to work so I can show what I'm exactly doing

OpenStudy (mendicant_bias):

(Nope, still not able to get it to work, but thanks for the moment, just trying to figure out why I can't get the LaTeX editor working/)

ganeshie8 (ganeshie8):

its not working from two days

OpenStudy (mendicant_bias):

This is happening to everybody?

ganeshie8 (ganeshie8):

nope, oly for some priviliged users...

OpenStudy (mendicant_bias):

Will do drawing tool. Ah.

OpenStudy (mendicant_bias):

Looks like I'm not able to use the drawing tool, either, lol...

ganeshie8 (ganeshie8):

yes both drawing+latex became buggy from two days for many users

hartnn (hartnn):

\(\Large \dfrac \int \sqrt{x^3-3}{x^{11}}\)

ganeshie8 (ganeshie8):

http://www.twiddla.com/1465496 cme here

OpenStudy (mendicant_bias):

Lol I actually opened TeX Studio and started manually typing it up, but twiddla is fine